# Graph Neural Network Surrogates to leverage Mechanistic Expert Knowledge towards Reliable and Immediate Pandemic Response

### v3

@@ Abstract
During the COVID-19 crisis, mechanistic models have guided evidence-based decision making. However, time-critical decisions in a dynamical environment limit the time available to gather supporting evidence. We address this bottleneck by developing a graph neural network (GNN) surrogate of a spatially and demographically resolved mechanistic metapopulation simulator. This combined approach advances classical machine learning approaches which are often black box. Our design of experiments spans outbreak and persistent-threat regimes, up to three contact change points, and age-structured contact matrices on a 400-node spatial graph. We benchmark multiple GNN layers and identify an ARMAConv-based architecture that offers a strong accuracy–runtime trade-off. Across 30–90 day horizons and up to three contact change points, the surrogate attains 10–27% mean absolute percentage error (MAPE) while delivering (near) constant runtime with respect to the forecast horizon. Our approach accelerates evaluation by up to 28 670 times compared with the mechanistic model, allowing responsive decision support in time-critical scenarios and straightforward web integration. These results show how GNN surrogates can translate complex metapopulation models into immediate, reliable tools for pandemic response.
@@
journal: Simulation Modelling Practice and Theory
@@
{graphicalabstract}[Uncaptioned image]
{highlights}
GNN surrogate matches metapopulation expert model at 10–27% MAPE
Up to 28 670 times faster; less than 0.1s0.1\,s per 100 predictions
[…]
Interventions modeled through three different contact change points
Integration of mobility adjacency matrix
@@ 1 Introduction
Mathematical models can help public health experts and decision makers to explore potential future outcomes of ongoing disease dynamics. In particular during the recent COVID-19 crisis, mathematical modeling has been one of the principal forms to provide evidence on the effectiveness of public health and social interventions [1]. With an estimated number of more than 600 000 undiscovered viruses in mammal and avian hosts that are capable of infecting humans [2], pandemic preparedness is a necessity.
@@
Due to local outbreaks and human contact patterns, infectious disease dynamics are often heterogeneous on a spatial or demographic scale. For efficient mitigation of pandemics and accurate predictions, local transmission dynamics should be considered and corresponding data should be integrated in a mathematical model. Over the last years, a large number of authors have made contributions to predict the development of SARS-CoV-2. Simple ordinary differential equation (ODE)-based SIR (Susceptible-Infected-Removed)-type models [3, 4] have been used for their efficiency in time-critical moments. For using more realistic, nonexponential disease stage transitions, linear chain trick, generalized linear chain trick, or flexible integro-differential equation-based models are available; see, e.g., [5, 6, 7, 8, 9].
@@
Through metapopulation modeling, various authors addressed the spatially heterogeneous spread of SARS-CoV-2 and modeled entire countries with high efficiency and small execution times [10, 11, 12, 13, 14, 15]. To include individual transmission and superspreading events in an intuitive way, different contact network-, agent-, or individual-based models have been developed; see, e.g., [16, 17, 18, 19]. However, while more complex models allow more detailed research questions to be answered, they generally come at higher computational costs. Recent works [20, 21, 22] combine agent- and equation-based metapopulation approaches to substantially reduce the prediction time of complex agent-based models, yet these models remain the most computationally demanding among the involved approaches. Furthermore, many models using artificial intelligence (AI) have been proposed to fight the SARS-CoV-2 pandemic. While these models often deliver accurate results in short execution times, a review on more than 650 AI-based models identified most of them to be black box models [23]. Unfortunately, black box models often lack interpretability or explainability and do not contribute to a better understanding of underlying processes—important when deciding and communicating interventions to the public. Nevertheless, time is a critical factor in pandemics, and evaluations of mitigation and reaction strategies have to be conducted in narrow time windows. In addition, interventions ranging from distancing recommendations to face masks (in particular locations), capacity-restrictions, and closures of venues or even vaccinations (if available) yield a tremendously large combinatorial space from which decision makers need to select. While very simplified models are executed within the range of (milli)seconds, complex models, even if highly optimized, might need seconds to minutes to be run. With automated pipelines, preventing manual delay, and high-performance computing [24], thousands or millions of outcomes can be computed on a supercomputing facility to inform decision makers. However, in a dynamically changing environment, decision makers might need expected outcomes for changing initial conditions or other combinations of nonpharmaceutical interventions (NPIs). In addition to NPIs, public recommendations or voluntary behavioral change that can lead to new contact patterns have to be represented in a mathematical model. Eventually, low-barrier access to compute outcomes on all these scenarios through, e.g., a web application, can drastically increase the impact of the developed models.
@@
In this paper, we combine an already validated spatially and demographically resolved expert mechanistic model with AI-based approaches to provide an on-the-fly execution of spatially resolved infectious disease models. In [25], the authors compared advantages and disadvantages of several established machine learning approaches such as random forests, decision trees, and neural networks as surrogates for agent-based models (ABMs). Recently, the authors of [26] employed random forests to calibrate a city-scaled ABM. Here, however, we build upon an already validated spatially and demographically resolved metapopulation model [27], whose simulations are used to train a machine learning model with spatial resolution.
@@
The model proposed by [27] interprets the metapopulation model as a graph, with nodes representing subpopulations and edges capturing the exchange between them. With the expert model executed millions of times on a supercomputing facility overnight, we obtain a diverse training set from which a neural network surrogate can learn and generalize. As the output of the model from [27] is already in the form of a graph, a graph-based machine learning model is a natural choice and we therefore propose to employ graph neural networks (GNNs). GNNs are powerful tools for spatially distributed data, which have been introduced in [28] and further popularized in [29]. While we propose GNNs as surrogates for expert mechanistic models of infectious disease dynamics, GNNs have already excelled on data in different applications such as weather predictions [30], influenza-like diseases [31], or even SARS-CoV-2 [32, 33, 34].
@@ 2.1 Mechanistic model
The underlying mechanistic expert model which generates the reference training has to integrate the most important disease states and properties of disease dynamics for the surrogate model to learn from. This is required for obtaining a surrogate model that is reliable in practice. While deaths are usually monitored, several indicators such as the number of reported cases, the number of hospitalizations or intensive care unit (ICU) admissions have been considered in Germany to implement or lift nonpharmaceutical interventions against SARS-CoV-2 [35]. Consequently, we have chosen an epidemic model based on ordinary differential equations (ODEs) that has already been validated in [27] and which categorizes subpopulations based on the infection states susceptible, exposed (not infectious), a- or presymptomatic and infectious, mildly symptomatic and infectious, severely infected (hospitalized), critically infected (ICU), dead, and recovered; see Fig. 1. Aside the mentioned indicators, the model particularly considers pre- and asymptomatic transmission as studies on SARS-CoV-2 observed that viral load is peaking close to symptom onset [36, 37]. The model is efficiently implemented in C++ in the MEmilio [38, 39] framework, which means that it yields very fast model evaluations on a single core. We furthermore divided the population into $G=6$ age groups, that is, from 0 to 4, 5 to 14, 15 to 34, 35 to 59, 60 to 79, and 80+ years.
@@
$\frac{dz_{j,i}(t)}{dt}$ $=\mu_{z_{j-1,i}}^{z_{j,i}}\frac{z_{j-1,i}(t)}{T_{z_{j-1,i}}}-\frac{z_{j,i}(t)}{T_{z_{j,i}}},$
where $z_{j-1,i}$ is the former, less severe state. For the full set of equations, we refer to [27].
In this study, we use parameters that were obtained for the first waves of wild-type SARS-CoV-2 in 2020. Identical to [6], the parameters are either directly obtained or calculated based on [27, Table 2]. For the average values over all age groups, the age-specific values are weighted with respect to share of the corresponding age group in the total population provided for Germany by [40]. The overview on the particular values can be found in Tables 1 and 2.
@@
Table 2: Age-resolved parameters for wild-type SARS-CoV-2 based on [27].
@@
where $\phi_{i,j,0}$ is the initial contact rate, $\delta$ defines a transition interval, and $\widehat{\phi}_{i,j}$ represents a transition contact rate to ensure smooth transition, i.e., ${\phi}_{i,j}(t)\in\mathcal{C}^{1}((c_{1}-\varepsilon,c_{1}+\delta+\varepsilon))$ for ${i,j}\in\{1,\ldots,6\}$ with $0<\varepsilon\ll 1$. As suggested in [27], $\widehat{\phi}$ is realized through a cosine function.
In order to realize the spatial resolution, we employed a generalized graph as described in [27]. Proceeding with this approach, we obtained a graph with one node for each spatial region, which is then equipped with a a local disease dynamics model based on ODEs. Furthermore, we obtained one edge for each pair of age group and disease state, which then implements the corresponding mobility exchange. The $400\times 400$ mobility matrix for all of Germany’s counties was extrapolated from the German Federal Employment Agency [41]. The adjacency matrix $A$ needed for the GNN was obtained from this mobility data matrix by setting all nonzero connections to one. While also this metapopulation model was implemented efficiently and did undergo several optimization steps, for large experiments we still require relevant computing resources; cf. Fig. 6, Panel C.
@@ 2.2 Performance evaluation and objective function
From our application, we obtain time-dependent (and spatially resolved) outputs. The output vectors from all models are interpolated to full simulation days so that corresponding demographically and spatially resolved data sets have the same dimension. To evaluate the prediction performance of the neural network-based surrogate models as well as the dummy estimator (see next subsection), we use the mean absolute percentage error (MAPE) as evaluation metric, which is defined as
@@
as objective function. Here, $n$ is the number of samples, $y_{i}$ are the interpolated ”ground truth” daily values and $\widehat{y}_{i}$ are the predicted daily values.
@@ 2.4 Neural network surrogate model
We conduct a multi-stage model building and optimization process to build up a neural network with spatially resolved infection predictions. Before deriving suitable input data structures for learning a spatially resolved graph-based model with a GNN, we consider the temporal evolution of disease dynamics. For this, we train an MLP, an LSTM and a CNN model, each implemented using the Keras [42] library with TensorFlow [43] as backend.
@@
After our preliminary tests with simple neural network models which do not consider spatial effects, we turn to GNNs. We use the Spektral library [44] for defining the graph data and building the GNNs and Keras to train and evaluate the model. Spektral is an open-source Python library based on the Keras API and TensorFlow 2 [43] implementing GNN models. Other existing libraries, which are not considered here, are PyTorch Geometric [45] and DeepGraphLibrary [46].
GNNs can be categorized into spectral and spatial approaches, where spectral approaches have a mathematical foundation in graph signal processing while spatial approaches directly aggregate information from the nodes’ spatial neighborhoods [47], which is given by the graph. We compare the performance of four different types of GNN layers available in Spektral to build our model. As a first approach, we consider GCNConv layers, which are based on the graph convolutional network (GCN) as proposed by [48]. In this spectral approach, a modified adjacency matrix with self-loops is used. The second type are ARMAConv layers, which are based on [49]. Like the GCNConv layer, this is a spectral approach, and it applies convolutional operations in combination with an autoregressive moving average (ARMA) approach for improved training performance; cf. the discussion below. The third approach uses approximative personalized propagation of neural predictions convolutional (APPNPConv) layers based on [50], which can also be categorized as a spectral approach with convolutional operations. Lastly, we consider the graph attention convolutional (GATConv) layer from [51], which is different from the other considered approaches, in the sense that it uses a spatial approach that implements an attention mechanism.
Since the ARMAConv layer performed best (see the Results section), we briefly provide some details on this specific type of GNN layer as well as our final architecture. In [49], the authors substituted polynomial filters in GNNs by filters which are based on the ARMA approach. In particular, in our optimized architecture, the final model consists of seven ARMAConv layers with 512 channels per layer and consisting of one stack of GCS layers, which serves as a filter on the node feature input data. Generally, we denote a ARMAConv layer with $k$ GCS stacks as $\textit{ARMA}_{k}$. The $i$th GCS layer in the stack is parametrized by the trainable weight matrix $\textbf{W}_{i}$ as well as the trainable skip connection $\textbf{V}_{i}$, which arises from the ARMA scheme applied to the GCS stack. In the case of the $\textit{ARMA}_{1}$ layer with one GCS stack, the graph convolution layer corresponds to

### v4

@@ Abstract
During the COVID-19 crisis, mechanistic models have guided evidence-based decision making. However, time-critical decisions in a dynamical environment limit the time available to gather supporting evidence. We address this bottleneck by developing a graph neural network (GNN) surrogate of an age-structured and spatially resolved mechanistic metapopulation simulation model. This combined approach complements classical modeling approaches which are mostly mechanistic and purely data-driven machine learning approaches which are often black box. Our design of experiments spans outbreak and persistent-threat regimes, up to three contact change points, and age-structured contact matrices on a spatial graph with 400 nodes representing German counties. We benchmark multiple GNN layers and identify an ARMAConv-based architecture that offers a strong accuracy–runtime trade-off. Across horizons of 30–90 day simulation and prediction, allowing up to three contact change points, the surrogate model attains 10–27% mean absolute percentage error (MAPE) while delivering (near) constant runtime with respect to the forecast horizon. Our approach accelerates evaluation by up to 28 670 times compared with the mechanistic model, allowing responsive decision support in time-critical scenarios and straightforward web integration. These results show how GNN surrogates can translate complex metapopulation models into immediate, reliable tools for pandemic response.
@@
firstfirstfootnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Mathematical models can help public health experts and decision makers to explore potential future outcomes of ongoing disease dynamics. In particular during the recent COVID-19 crisis, mathematical modeling has been one of the principal forms to provide evidence on the effectiveness of public health and social interventions ecdc_2024 . With an estimated number of more than 600 000 undiscovered viruses in mammal and avian hosts that are capable of infecting humans daszak_workshop_2020 , pandemic preparedness is a necessity.
@@
Due to local outbreaks and human contact patterns, infectious disease dynamics are often heterogeneous on a spatial or demographic scale. For efficient mitigation of pandemics and accurate predictions, local transmission dynamics should be considered and corresponding data should be integrated in a mathematical model. Over the last years, a large number of authors have made contributions to predict the development of SARS-CoV-2. Simple ordinary differential equation (ODE)-based SIR (Susceptible-Infected-Removed)-type models bauer_relaxing_2021 , schuler_data_2021 have been used for their efficiency in time-critical moments. For using more realistic, nonexponential disease stage transitions, linear chain trick (see, e.g., macdonald_time_1978 , PLOTZKE2026823 ), generalized linear chain trick hurtado_generalizations_2019 , or flexible integro-differential (see, e.g., medlock_spreading_2003 , WENDLER2026129636 ) equation-based models are available.
@@
Through metapopulation modeling, various authors addressed the spatially heterogeneous spread of SARS-CoV-2 and modeled entire countries with high efficiency and small execution times pei_differential_2020 , chen_compliance_2021 , levin_effects_2021 , liu_modelling_2022 , zunker_novel_2024 , ZUNKER2025116782 . For instance, Pei et al. pei_differential_2020 considered effects of intervention timing on COVID-19 spread in the United States (US), Chen et al. chen_compliance_2021 considered compliance and containment of COVID-19 across towns in the US, Levin et al. levin_effects_2021 considered effects of short-term travel on COVID-19 spread in Minnesota, while Liu et al. liu_modelling_2022 considered COVID-19 cases in Singapore and Zunker et al. zunker_novel_2024 , ZUNKER2025116782 considered travel-time aware and behavior-driven metapopulation models with the 400 German counties. To include individual transmission and superspreading events in an intuitive way, different contact network-, agent-, or individual-based models have been developed; see, e.g., bershteyn_implementation_2018 , kerr_covasim_2021 , muller_predicting_2021 , KERKMANN2025110269 . For instance, Bershteyn et al. provided a very general pre-COVID-19 agent-based model (ABM) with different transmission pathways (vector, environmental, or airborne), Kerr et al. kerr_covasim_2021 developed a powerful ABM for COVID-19, Müller et al. muller_predicting_2021 extended the MATSim ABM for the COVID-19 use case, and Kerkmann et al. KERKMANN2025110269 developed an ABM for respiratory diseases and direct or airborne transmission pathways. However, while more complex models allow more detailed research questions to be answered, they generally come at higher computational costs. Recent works with very similar approaches bicker_hybrid_2024 , kehrer_hybrid_2025 , bostanci2024 combine agent- and equation-based metapopulation approaches to substantially reduce the prediction time of complex ABMs, yet these models remain the most computationally demanding among the involved approaches. Furthermore, many models using artificial intelligence (AI) have been proposed to fight the SARS-CoV-2 pandemic; see, e.g., tayarani_n_applications_2021 . While these models often deliver accurate results in short execution times, a review on more than 650 AI-based models identified most of them to be black box models tayarani_n_applications_2021 . Unfortunately, black box models often lack interpretability or explainability and do not contribute to a better understanding of underlying processes—important when deciding and communicating interventions to the public. Nevertheless, time is a critical factor in pandemics, and evaluations of mitigation and reaction strategies have to be conducted in narrow time windows. In addition, interventions ranging from distancing recommendations to face masks (in particular locations), capacity-restrictions, and closures of venues or even vaccinations (if available) yield a tremendously large combinatorial space from which decision makers need to select. While very simplified models are executed within the range of (milli)seconds, complex models, even if highly optimized, might need seconds to minutes to be run. With automated pipelines preventing manual delay and high-performance computing memon_automated_2024 , thousands or millions of outcomes can be computed on a supercomputing facility to inform decision makers on short term. However, in a dynamically changing environment, decision makers might need expected outcomes for individual combinations of nonpharmaceutical interventions (NPIs) on-the-fly. Training a GNN on the HPC simulations eventually allows the user-driven exploration of any strictness of NPIs, based on the interpolation capabilities of the GNN. Through a low-barrier access for public health experts, e.g., a web application, this approach also drastically increases the impact of the developed (mechanistic) models. Furthermore, novel NPIs, public recommendations, or observed behavioral changes will be represented in the mechanistic models with new contact patterns and learned regularly with new simulation outputs by novel versions of the GNN.
@@
In this paper, we combine an already validated age-structured and spatially resolved expert mechanistic model kuhn_assessment_2021 , koslow_appropriate_2022 with AI-based approaches to provide an on-the-fly execution of spatially resolved infectious disease models. A schematic presentation of the approach is shown in Fig. 1. In angione_using_2022 , Angione et al. compared advantages and disadvantages of several established machine learning approaches such as random forests, decision trees, and neural networks as surrogates for agent-based models. Recently, Robertson et al. robertson_bayesian_2024 employed random forests to calibrate a city-scaled ABM. Here, however, we build upon an already validated age-structured and spatially resolved metapopulation model kuhn_assessment_2021 , whose simulations are used to train a machine learning model with spatial resolution.
Figure 1: Visual representation of the prediction phase for a GNN surrogate model. The adjacency matrix results from the links between the considered regions, here depicted high-level as European countries. The magnifying glass highlights mobility connections between Switzerland and other countries. Colors on the national level geography represent disease outcomes such as symptomatic cases incidence, hospitalization, or deaths. Visualization of a GNN with four ARMAConv layers with three Graph Convolutional Skip (GCS) stacks and one dense output layer. Visualization based on Eurostat shape files eurostat © EuroGeographics for the administrative boundaries.
@@
The model proposed by Kühn et al. kuhn_assessment_2021 interprets the metapopulation model as a graph, with nodes representing subpopulations and edges capturing the exchange between them. With the expert model executed millions of times on a supercomputing facility overnight, we obtain a diverse training set from which a neural network surrogate can learn and generalize. As the output of the model from kuhn_assessment_2021 is already in the form of a graph, a graph-based machine learning model is a natural choice and we therefore propose to employ graph neural networks (GNNs). GNNs are powerful tools for spatially distributed data, which have been introduced in Gori et al. gori_new_2005 and further popularized in Scarselli et al. scarselli_graph_2009 . While we propose GNNs as surrogates for expert mechanistic models of infectious disease dynamics, GNNs have already excelled on data in different applications such as weather predictions Lam2023 , influenza-like diseases deng_graph_2019 , or even SARS-CoV-2 kapoor_examining_2020 , gao_stan_2021 , panagopoulos_transfer_2021 , Fritz2022 . Fritz et al. Fritz2022 used GNNs in a purely data-driven fashion for spatio‐temporal disease models, Gao et al. gao_stan_2021 used patients’ claims data from different counties in a graph-based approach with a transmission dynamics loss, and Panagopoulos et al. panagopoulos_transfer_2021 used transfer learning together with reported case data-driven GNNs.
@@ 2.1 Mechanistic model
A requirement for obtaining a surrogate model that is reliable in practice is that the model, which is replaced by the surrogate, has been validated beforehand. For infectious disease dynamics, the underlying mechanistic expert model that generates the reference training has to integrate the most important disease states and properties of disease dynamics for the surrogate model to learn from. While deaths are usually monitored, several indicators such as the number of reported cases, the number of hospitalizations or intensive care unit (ICU) admissions have been considered in Germany to implement or lift nonpharmaceutical interventions against SARS-CoV-2 rki_control_covid_2021 . Consequently, we have chosen an epidemic model based on ordinary differential equations (ODEs) that has already been validated in kuhn_assessment_2021 , koslow_appropriate_2022 and which categorizes subpopulations based on the infection states susceptible, exposed (not infectious), a- or presymptomatic and infectious, mildly symptomatic and infectious, severely infected (hospitalized), critically infected (ICU), dead, and recovered; see Fig. 2. Aside the mentioned indicators, the model particularly considers pre- and asymptomatic transmission as studies on SARS-CoV-2 observed that viral load is peaking close to symptom onset as observed by Walsh et al. and Puhach et al. walsh_duration_2020 , puhach_sars-cov-2_2023 . The model is efficiently implemented in C++ in the MEmilio Bicker_MEmilio_2026 framework, which means that it yields very fast model evaluations on a single core. We furthermore divided the population into $G=6$ age groups, that is, from 0 to 4, 5 to 14, 15 to 34, 35 to 59, 60 to 79, and 80+ years. While also simpler models of SIR-type structure and without age-structure could be learnt by a surrogate model, the resulting surrogate model could then not be evaluated for critical numbers on hospitalizations, intensive care, or deaths and without learning about most contributing age groups. In addition, the SIR-type structure would miss the well observed disease evolution of COVID-19 with exposed, non-infectious and nonsymptomatic, infectious states.
@@
$\frac{dz_{j,i}(t)}{dt}$ $=\mu_{z_{j-1,i}}^{z_{j,i}}\frac{z_{j-1,i}(t)}{T_{z_{j-1,i}}}-\frac{z_{j,i}(t)}{T_{z_{j,i}}},$ (2)
where $z_{j-1,i}$ is the former, less severe state. For the full set of equations, we refer to kuhn_assessment_2021 .
In this study, we use parameters that were obtained for the first waves of wild-type SARS-CoV-2 in 2020. Identical to Plötzke et al. PLOTZKE2026823 , the parameters are either directly obtained or calculated based on [kuhn_assessment_2021, , Table 2]. For the average values over all age groups, the age-specific values are weighted with respect to share of the corresponding age group in the total population provided for Germany by regionaldatenbank_deutschland_fortschreibung . The overview on the particular values can be found in Tables 1 and 2.
@@
Table 2: Age-resolved parameters for wild-type SARS-CoV-2 based on kuhn_assessment_2021 .
@@
where $\phi_{i,j,0}$ is the initial contact rate, $\delta$ defines a transition interval, and $\widehat{\phi}_{i,j}$ represents a transition contact rate to ensure smooth transition, i.e., ${\phi}_{i,j}(t)\in\mathcal{C}^{1}((c_{1}-\varepsilon,c_{1}+\delta+\varepsilon))$ for ${i,j}\in\{1,\ldots,6\}$ with $0<\varepsilon\ll 1$. As suggested in kuhn_assessment_2021 , $\widehat{\phi}$ is realized through a cosine function.
In order to realize the spatial resolution, we employed a generalized graph as described in kuhn_assessment_2021 . Proceeding with this approach, we obtained a graph with one node for each spatial region, which is then equipped with a a local disease dynamics model based on ODEs. Furthermore, we obtained one edge for each pair of age group and disease state, which then implements the corresponding mobility exchange. The $400\times 400$ mobility matrix for all of Germany’s counties was extrapolated from the German Federal Employment Agency bmas_pendlerverflechtungen_2020 . The adjacency matrix $A$ needed for the GNN was obtained from this mobility data matrix by setting all nonzero connections to one.
The resulting graph, representing 400 German counties, is directed and has a density of $30.4\%$ with $48,467$ edges. Its topology is characterized by an uneven degree distribution with highly connected hubs (e.g., Frankfurt) connecting to smaller surrounding areas, indicated by a slightly negative assortativity coefficient (0.096-0.096). A high clustering coefficient ($0.56$) further reflects strong local community structures. From a physical perspective, the GNN learns to approximate the epidemic diffusion process on this specific topology. Since the local rules of this diffusion are consistent, the model is capable of generalizing to unseen temporal scenarios, given the local connectivity patterns remain similar.
While also this metapopulation model was implemented efficiently and did undergo several optimization steps, for large experiments we still require relevant computing resources; cf. Fig. 9, Panel C.
@@ 2.2 Performance evaluation and objective function
Simulating the previously described mechanistic metapopulation model, we obtain time-dependent (and spatially resolved) outputs. The output vectors from all local models of the form Eqs. 1 and 2 are interpolated to full simulation days so that corresponding age-structured and spatially resolved data sets have the same dimension. To evaluate the prediction performance of the neural network-based surrogate models as well as the dummy estimator (see next subsection), we use the mean absolute percentage error (MAPE) as evaluation metric, which is defined as
@@
as objective function. Here, $n$ is the total number of data points which is the product of the number of runs, the number of days, and the number of compartments, $y_{i}$ and $\widehat{y}_{i}$ denote individual scalar values representing the interpolated ”ground truth” and predicted daily values, respectively
@@ 2.4 Neural network surrogate model
We conduct a multi-stage model building and optimization process to build up a neural network with spatially resolved infection predictions. Before deriving suitable input data structures for learning a spatially resolved graph-based model with a GNN, we consider the temporal evolution of disease dynamics. For this, we train an MLP, an LSTM and a CNN model, each implemented using the Keras chollet_keras_2015 library with TensorFlow tensorflow2015 as backend.
@@
After our preliminary tests with simple neural network models which do not consider spatial effects, we turn to GNNs. We use the Spektral library from Grattarola and Alippi grattarola_graph_2021 for defining the graph data and building the GNNs and Keras to train and evaluate the model. Spektral is an open-source Python library based on the Keras API and TensorFlow 2 tensorflow2015 implementing GNN models. Other existing libraries, which are not considered here, are PyTorch Geometric fey_fast_2019 and DeepGraphLibrary wang_deep_2020 .
GNNs can be categorized into spectral and spatial approaches, where spectral approaches have a mathematical foundation in graph signal processing while spatial approaches directly aggregate information from the nodes’ spatial neighborhoods (see Wu et al. wu_comprehensive_2021 ), which is given by the graph. We compare the performance of four different types of GNN layers available in Spektral to build our model. As a first approach, we consider GCNConv layers, which are based on the graph convolutional network (GCN) as proposed by Kipf et al. kipf_semi-supervised_2017 . In this spectral approach, a modified adjacency matrix with self-loops is used. The second type are ARMAConv layers, which are based on Bianchi et al. bianchi_graph_2021 . Like the GCNConv layer, this is a spectral approach, and it applies convolutional operations in combination with an autoregressive moving average (ARMA) approach for improved training performance; cf. the discussion below. The third approach uses approximative personalized propagation of neural predictions convolutional (APPNPConv) layers based on Gasteiger et al. gasteiger_predict_2022 , which can also be categorized as a spectral approach with convolutional operations. Lastly, we consider the graph attention convolutional (GATConv) layer from Velickovic et al. velickovic_graph_2018 , which is different from the other considered approaches, in the sense that it uses a spatial approach that implements an attention mechanism.
Since the ARMAConv layer performed best (see the Results section), we briefly provide some details on this specific type of GNN layer as well as our final architecture. In bianchi_graph_2021 , Bianchi et al. substituted polynomial filters in GNNs by filters which are based on the ARMA approach. In particular, in our optimized architecture, the final model consists of seven ARMAConv layers with 512 channels per layer and consisting of one stack of GCS layers, which serves as a filter on the node feature input data. Generally, we denote a ARMAConv layer with $k$ GCS stacks as $\textit{ARMA}_{k}$. The $i$th GCS layer in the stack is parametrized by the trainable weight matrix $\textbf{W}_{i}$ as well as the trainable skip connection $\textbf{V}_{i}$, which arises from the ARMA scheme applied to the GCS stack. In the case of the $\textit{ARMA}_{1}$ layer with one GCS stack, the graph convolution layer corresponds to